OpenStudy (anonymous):

What is the value of the expression? 7^-3x7^9/7^-8 A. 714 B. 72 C. 7−14 D. 7−2

OpenStudy (johnweldon1993):

Come on you got this :P \[\large \frac{7^{-3}\times 7^{9}}{7^{-8}}\] Take care of the top first...leave the 7 alone and add the exponents...so -3 + 9 = 6 So now we have \[\large \frac{7^6}{7^{-8}}\] And now we leave the 7 alone again and subtract the exponents \[\large 6 - (-8) = ?\]

OpenStudy (anonymous):

so its -14 right?

OpenStudy (johnweldon1993):

Nope... \[\large 6 - (-8) = 6 + 8 = ?\]

OpenStudy (anonymous):

oh 14

OpenStudy (anonymous):

srry

OpenStudy (johnweldon1993):

Lol dont worry about it good job :P